1) Betsy, a recent retiree, requires $6000 per year in extra income. She has $50,000 to invest and can in­ vest in B-rated bonds paying 15% per year or in a certificate of deposit (CD) paying 7% per year. How much money should be invested in each to realize exactly $6000 in inter­est per year?
2) After 2 years, Betsy (see Problem 1) finds that she will now require $7000 per year. Assuming that the remaining information is the same, how should the money be reinvested?

Answer by ewatrrr(24785)


Hi 
Let 'x' represent amount invested at 15%
I. .07(50000-x) + .15x = $6000
   .08x = 6000-.07*50,000
      x = 2500/.08
      x =$31,250  invested at 15%, $18,750 at 7%
II.
Set up the same:
Let 'x' represent amount invested at 15%
 .07(50000-x) + .15x = $7000
  .08x = 7000 - 3500
     x = 3500/.08
     x = $43,750 at 15%,  $6,250 at 7%

Wish You the Best in your Studies.